A detector of radioactivity in a laboratory indicates an average of 16 counts /min when no radioactive samples are present. A radioactive sample of half-life 1.5 days is placed close to the detector, which indicates a count rate of 208 counts /min. Calculate the count rate that is indicated 6 days later.
1
Expert's answer
2020-06-03T12:04:44-0400

actual initial count = 208-16=192 counts /min

as we know that after every half life counts gets reduced to half

after 1.5 days count= 192/2 = 96 counts /min

after next 1.5 days(total = 3 days) = 96/2 = 48 counts /min

after next 1.5 days (total = 4.5 days)= 48/2 = 24 counts / min

after next 1.5 days (total = 6 days) =24/2 = 12 counts/min


As we are given there is average count in the lab without any sample hence we will add 16 counts to our answer

Final count = 12+16 = 28 count/min